the sail on a sailboat is in the shape of a triangle. if the height of the sail is 5 times the length of base and the area is 160 square feet, find the dimensions of the sail.

Answer by texttutoring(324)

Area of a triangle is A=b*h/2

You know that the height is 5 times the length of the base:
h = 5b

You also know that the Area is 160 square feet. Put all this info together:

A=b*h/2
160 = b*5b / 2
160 = 5b^2 / 2
320 = 5b^2
64 = b^2
b = 8

So this means the base is 8 feet.

The height is 5 times the base, so it is 40 feet.